a tank can be filled in 45 minutes and by another pipe in 55 minutes. how long will it take both to fill it?
Answer by ewatrrr(10682)


Hi
Let x represent time in minutes it takes to fill the tank together
Per minute is the Equalizer
1/45 + 1/55 = 1/x Mulitplying each term by 45*55x so as all denominators = 1
55x + 45x = 45*55
100x = 45*55
x = 45*55/100
x = 27.75 min, time in minutes it takes to fill the tank together